# Gravitational waves and black holes from the phase transition in models of dynamical symmetry breaking

### v2

@@ 1 Introduction
Among various cosmological mechanisms for producing a SGWB [26], cosmological strong first-order phase transitions (FOPTs) (see e.g. [27]) offer plausible beyond the Standard Model (BSM) signals, up to high scales. Indeed the SM predicts two (electroweak and QCD) phase transitions, none of which is of first-order [28, 29]. Therefore, the detection of a GW signal compatible with a FOPT would be evidence of BSM physics. FOPTs develop by the formation of bubbles that expand, collide and percolate. The violent collisions between the bubble walls (and the motion of the surrounding thermal plasma) lead to the production of stochastic GWs. The typical frequency of such gravitational waves is the red-shifted Hubble rate during the phase transition, $f_{\rm peak}\sim T_{0}T/M_{\rm Pl}$. A FOPT around the electroweak scale, $T\sim 100$ GeV, would peak at $f\sim$ mHz [30] which is in the frequency sensitivity band of space-based GW experiments such as LISA [15], whereas ground-based experiments such as LIGO-VIRGO [31, 32] and ET [20] around 100 Hz can probe FOPTs up to $T\sim 10^{8}$ GeV [33], beyond the reach of collider experiments.
@@
In this work we analyze the cosmological implications of such BSM phase transitions. Besides an enhanced GW signal, a supercooled FOPT also leads to the formation of primordial black holes (PBHs) via different mechanisms. Bubble collisions have been studied e.g. in [58]. We focus on a mechanism suggested in [59, 60, 61, 62, 63, 64] that recently received interest [65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87]. FOPTs proceed via the nucleation of bubbles of the broken phase in an initial background of the symmetric phase [88, 89, 90]. In a supercooled regime, the energy density of the Universe in the symmetric phase is dominated by the vacuum energy which acts as a cosmological constant and leads to an inflationary period. In a nucleated bubble, instead, such energy is quickly converted into radiation and the corresponding patch expands slower. Since bubble nucleation is a stochastic phenomenon, and since, in a supercooled regime, regions outside the nucleated bubbles dilute faster than those inside, a region where nucleation happens later becomes over-dense and eventually collapses, if large enough, into a PBH.
@@ 2.2 Computing the phase transition
so that the phase transition happens in $H/\beta$ Hubble times. The duration of the phase transition gets longer, up to a fraction of Hubble time, in the regime where $T_{\rm nuc}$ is much smaller than the mass scales of the problem. This happens because $d(S_{3}/T)/d\ln T\propto\ln^{-2}(w/T)$ decreases at $T\ll w$, where the approximation of eq. (2.11) gets more accurate. As discussed later in section 2.4, a substantial amount of primordial black holes forms in this regime, near to the critical boundary where nucleation becomes too slow for ending super-cooling. If $T$ gets smaller than $\Lambda_{\rm QCD}$, super-cooling is ended by QCD effects [36, 43].
@@ 2.4 PBH formation from strong first-order phase transition
A first order phase transition can lead to formation of black holes [68, 73, 80, 79] in different ways. In particular, due to the stochastic nature of nucleation, there is a probability that some regions remain for a longer time in the false vacuum while the space around them gets filled by true vacuum. Re-heating in the true vacuum leads to radiation, so that its energy density $\rho_{\rm true}$ dilutes, while the energy density in the false vacuum is $\rho_{\text{false}}\simeq V_{\Lambda}$. Thus regions in the false vacuum become relatively denser as quantified by $\delta\equiv\rho_{\text{false}}/\rho_{\text{true}}-1$. If a region in the false vacuum (approximated as roughly-spherical) has radius larger than roughly the Hubble parameter, it forms a primordial black hole when the density contrast exceeds the critical value $\delta_{\text{c}}\approx 0.45$ [109].
@@
$M_{\text{PBH}}\approx k\frac{4\pi}{3}\left(\frac{c_{s}}{H}\right)^{3}\rho_{\rm rad}^{\rm late}\bigg{|}_{t_{\rm coll}}\approx M_{\odot}\left(\frac{20}{g_{\star}(T_{\text{eq}})}\right)^{1/2}\left(\frac{0.14\,\text{GeV}}{T_{\text{eq}}}\right)^{2}$ (2.28)
where $g_{\star}\approx 100$, $M_{\odot}\approx 2\times 10^{30}\,\text{kg}$ is the solar mass, $c_{s}\simeq 1/\sqrt{3}$ is the sound speed, $k\lesssim 1$, $\rho_{\rm rad}^{\rm late}$ is the false vacuum energy in a late-patch [80]. The DM range in eq. (2.27) roughly corresponds to equilibrium temperatures $50\,{\rm TeV}\lesssim T_{\rm eq}\lesssim 10\,{\rm PeV}$.
@@
The parameter $\gamma$ typically lies in the range $0.81-1$; we assume a reference value $\gamma=0.96$ [115, 116].
@@ 3 Minimal model with singlet scalar DM
$\lambda_{SS^{\prime}}$ $\approx\frac{(4\pi)^{2}m_{s}^{2}}{m_{s^{\prime}}^{2}},\qquad\lambda_{HS^{\prime}}$ $\approx-\frac{(4\pi)^{2}m_{h}^{2}}{m_{s^{\prime}}^{2}\ln R},\qquad w\simeq\frac{\sqrt{2}m_{s^{\prime}}}{4\pi m_{s}}.$ (3.4)
@@
with $f_{N}\approx 0.3$ and $m_{N}\approx 0.946\,{\rm GeV}$. The cosmological DM abundance is reproduced when $\sigma_{0}\approx 1/M^{2}$ with $M\approx 23\,{\rm TeV}$. Current bounds on direct detection are satisfied for $m_{s^{\prime}}\gtrsim 2\,{\rm TeV}$ [117, 120].
@@
Figure 1: Gravitational wave spectrum as a function of frequency $f$ at the benchmark points A, B, C, D listed in table 1. The predicted gravitational wave signals (colored curves) fall within the sensitivity curves of possible future gravitational wave detectors (dashed shaded regions). The upper axis shows the mass of the corresponding primordial black holes. The gray hatched curve shows the expected astrophysical foreground.
@@ 3.1 Primordial Black Holes as DM
The GW signals are above the expected astrophysical backgrounds, shown by the hatched curve in fig. 1 (see [132] for a list of references).
@@ 3.2 General parameter space of the model
Fig. 5 shows contour levels of the nucleation temperature and of the inverse duration $\beta/H$ of the phase transition. Interesting values are obtained in a region ranging from too large non-perturbative couplings down to too small couplings, where the phase transition is ended in a too rapid way by QCD rather than by nucleation [81, 164]. Fig. 5 shows, in the same plane, contour values of the PBH masses. Fig. 5 shows contour values of the PBH abundance. As it is exponentially sensitive to parameter values, the figures focuses on a small portion of the parameter space where the PBH abundance is around the DM abundance.
@@ Acknowledgement
We thank Luca Marzola for collaboration and discussion during the initial stage of the project, Marcos Flores and Tomer Volansky for discussions.

### v3

@@ 1 Introduction
Among various cosmological mechanisms for producing a SGWB [26], cosmological strong first-order phase transitions (FOPTs) (see e.g. [27]) offer plausible beyond the Standard Model (BSM) signals, up to high scales. Indeed the SM predicts two (electroweak and QCD) phase transitions, none of which is of first-order [28, 29]. Therefore, the detection of a GW signal compatible with a FOPT would be evidence of BSM physics. FOPTs develop by the formation of bubbles that expand, collide and percolate. The violent collisions between the bubble walls (and the motion of the surrounding thermal plasma) lead to the production of stochastic GWs. The typical frequency of such gravitational waves is the red-shifted Hubble rate during the phase transition, $f_{\rm peak}\sim T_{0}T/M_{\rm Pl}$, where $T_{0}$ is the current CMB temperature. A FOPT around the electroweak scale, $T\sim 100$ GeV, would peak at $f\sim$ mHz [30] which is in the frequency sensitivity band of space-based GW experiments such as LISA [15], whereas ground-based experiments such as LIGO-VIRGO [31, 32] and ET [20] around 100 Hz can probe FOPTs up to $T\sim 10^{8}$ GeV [33], beyond the reach of collider experiments.
@@
In this work we analyze the cosmological implications of such BSM phase transitions. Besides an enhanced GW signal, a supercooled FOPT also leads to the formation of primordial black holes (PBHs) via different mechanisms. Bubble collisions have been studied e.g. in [58]. We focus on a mechanism suggested in [59, 60, 61, 62, 63, 64] that recently received interest [65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 87]. FOPTs proceed via the nucleation of bubbles of the broken phase in an initial background of the symmetric phase [88, 89, 90]. In a supercooled regime, the energy density of the Universe in the symmetric phase is dominated by the vacuum energy which acts as a cosmological constant and leads to an inflationary period. In a nucleated bubble, instead, such energy is quickly converted into radiation and the corresponding patch expands slower. Since bubble nucleation is a stochastic phenomenon, and since, in a supercooled regime, regions outside the nucleated bubbles dilute faster than those inside, a region where nucleation happens later becomes over-dense and eventually collapses, if large enough, into a PBH.
@@ 2.2 Computing the phase transition
so that the phase transition happens in $H/\beta$ Hubble times. The duration of the phase transition gets longer, up to a fraction of Hubble time, in the regime where $T_{\rm nuc}$ is much smaller than the mass scales of the problem. This happens because $d(S_{3}/T)/d\ln T\propto\ln^{-2}(w/T)$ decreases at $T\ll w$, where the approximation of eq. (2.11) gets more accurate. [^3] We do not need to expand the decay rate density up to higher second order in Taylor series in the time difference from the nucleation time $t_{n}$, $\gamma(t)=\gamma(t_{\rm nuc})\text{exp}[\beta(t-t_{\rm nuc})+\beta_{2}(t-t_{\rm nuc})^{2}+\cdots],$ (2.17) because, in the considered phase transitions during super-cooling, the second order coefficient $\beta_{2}$ negligibly arises by expanding the logarithmic dependence of $S_{3}/T$ on $T/w\approx e^{-H(t-t_{\rm nuc})}T_{\rm nuc}/w\gg 1$. As discussed later in section 2.4, a substantial amount of primordial black holes forms in this regime, near to the critical boundary where nucleation becomes too slow for ending super-cooling. If $T$ gets smaller than $\Lambda_{\rm QCD}$, super-cooling is ended by QCD effects [36, 43].
@@ 2.4 PBH formation from strong first-order phase transition
A first order phase transition can lead to formation of black holes [68, 73, 80, 79] in different ways. In particular, due to the stochastic nature of nucleation, there is a probability that some regions remain for a longer time in the false vacuum while the space around them gets filled by true vacuum. Re-heating in the true vacuum leads to radiation, so that its energy density $\rho_{\rm true}$ dilutes, while the energy density in the false vacuum is $\rho_{\text{false}}\simeq V_{\Lambda}$. Thus regions in the false vacuum become relatively denser as quantified by $\delta\equiv\rho_{\text{false}}/\rho_{\text{true}}-1$. Following [80, 81] we assume that a primordial black hole forms when the density contrast exceeds the critical value $\delta_{\text{c}}\approx 0.45$ [109] in a region in the false vacuum (approximated as roughly-spherical) with radius larger than roughly the Hubble parameter. For small perturbations, this value for $\delta_{c}$ indeed coincides with the one at the horizon-crossing time, and it does not differ much from that even for the larger perturbations of interest for PBH formation [110]. It may however mildly vary depending on the perturbation profile [111]. We ignore this issue, since the power spectrum induced by phase transitions does not have sharp features.
@@
$M_{\text{PBH}}\approx F\frac{4\pi}{3}\left(\frac{c_{s}}{H}\right)^{3}\rho_{\rm rad}^{\rm late}\bigg{|}_{t_{\rm coll}}\approx M_{\odot}\left(\frac{20}{g_{\star}(T_{\text{eq}})}\right)^{1/2}\left(\frac{0.14\,\text{GeV}}{T_{\text{eq}}}\right)^{2}$ (2.29)
where $g_{\star}\approx 100$, $M_{\odot}\approx 2\times 10^{30}\,\text{kg}$ is the solar mass, $c_{s}\simeq 1/\sqrt{3}$ is the sound speed, and $F\lesssim 1$ is a numerical factor which depends on the details of the gravitational collapse [113]. $\rho_{\rm rad}^{\rm late}$ is the false vacuum energy in a late-patch [80]. The DM range in eq. (2.28) roughly corresponds to equilibrium temperatures $50\,{\rm TeV}\lesssim T_{\rm eq}\lesssim 10\,{\rm PeV}$.
@@
The parameter $\gamma$ typically lies in the range $0.81-1$; we assume a reference value $\gamma=0.96$ [118, 120, 119]. In general, $\gamma$ is inversely proportional to the spectral width $\sigma$ of power spectrum [84, 121], taking into account all scales. In our case, scales smaller than the horizon do not contribute to the creation of PBHs. This leads to a monotonically increasing functional dependence of the PBH abundance on $\sigma$. The assumed value of $\gamma$ should be realistic in the physically relevant domain, where the PBH abundance constraints are satisfied [84, 121]. In our scenario, the scalar avoids a phase where its quanta dominate the energy density: it quickly transfers its energy to the SM radiation, opening the phase when PBHs are formed.
@@ 3 Minimal model with singlet scalar DM
$\lambda_{SS^{\prime}}$ $\approx\frac{(4\pi)^{2}m_{s}^{2}}{m_{s^{\prime}}^{2}},\qquad\lambda_{HS^{\prime}}$ $\approx-\frac{(4\pi)^{2}m_{h}^{2}}{m_{s^{\prime}}^{2}\ln R},\qquad w\simeq\frac{\sqrt{2}m^{2}_{s^{\prime}}}{4\pi m_{s}}.$ (3.4)
@@
with $f_{N}\approx 0.3$ and $m_{N}\approx 0.946\,{\rm GeV}$. The cosmological DM abundance is reproduced when $\sigma_{0}\approx 1/M^{2}$ with $M\approx 23\,{\rm TeV}$. Current bounds on direct detection are satisfied for $m_{s^{\prime}}\gtrsim 2\,{\rm TeV}$ [122, 125]. Then indirect detection bounds are satisfied [112].
@@
Figure 1: Gravitational wave spectrum as a function of frequency $f$ at the benchmark points A, B, C, D listed in table 1. The predicted gravitational wave signals (colored curves) fall within the sensitivity curves of possible future gravitational wave detectors (dashed shaded regions). The upper axis shows the mass of the corresponding primordial black holes. The gray hatched curve indicates the expected but significantly uncertain astrophysical foreground.
@@ 3.1 Primordial Black Holes as DM
The GW signals are above the expected astrophysical backgrounds, shown by the hatched curve in fig. 1. This is obtained summing the different expected sources, and is subject to an uncertainty by about an order of magnitude. We have not taken into account that future observatories are expected to be able to partially subtract the astrophysical backgrounds. These issues have been studied in many papers. For a short summary and a list of references see e.g. [137].
@@ 3.2 General parameter space of the model
Fig. 5 shows contour levels of the nucleation temperature and of the inverse duration $\beta/H$ of the phase transition. Interesting values are obtained in a region ranging from too large non-perturbative couplings down to too small couplings, where the phase transition is ended in a too rapid way by QCD rather than by nucleation [81, 169]. Fig. 5 shows, in the same plane, contour values of the PBH masses. Fig. 5 shows contour values of the PBH abundance. As it is exponentially sensitive to parameter values, the figures focuses on a small portion of the parameter space where the PBH abundance is around the DM abundance. In this region, PBH can drive a small-scale clustering of particle DM, thereby enhancing indirect detection signals.
@@ Acknowledgement
We thank Luca Marzola for collaboration and discussion during the initial stage of the project, Marcos Flores, Marek Lewicki and Tomer Volansky for discussions.